a particle is trapped in a one-dimensional region of 0.1nm. With what precision could one reasonably hope to measure the momentum of the particle
Answer #1

We know that

\Delta x \Delta p=h/4 \pi

\Delta x =0.1 nm

So

\Delta p=h/4 \pi\Delta x

\Delta p=6.626*10^{-34}/4 \pi*10^{-10}

\Delta p=5.278*10^{-25} Kgm/s